### v1

# Revealing the Geometrical and Vibrational Properties of the Defects Driving the Boson Peak

Abstract
The vibrational density of states is key to understanding the mechanical, thermal, and transport properties of materials. In amorphous solids, this density shows an excess of vibrational modes compared to the Debye model, known as the boson peak, whose origin remains poorly understood. Previous studies have suggested a link between the boson peak and quasi-localized nonphononic vibrations, or "defects." However, it has been difficult to clearly identify these defects, possibly because they hybridize with extended phonons, casting doubt on their existence and connection to the boson peak. In this work, we introduce a simple and practical method for separating hybridized phonons from localized vibrations. We show that phonons at the boson peak frequency hybridize with localized defects. These defects are anisotropic, compact, and exhibit oscillatory pure shear deformations. Their density correlates with the excess of vibrational modes at the boson peak frequency across various two- and three-dimensional systems, confirming that they are the microscopic origin of the boson peak.
Thermal and vibrational properties of amorphous solids display universal characteristics that are very different from their crystalline counterparts [1]. In particular, their vibrational density of states (vDOS) $D(\omega)$ presents a universal excess anomaly over the Debye prediction $D(\omega)\propto\omega^{D-1}$, with $D$ the number of spatial dimensions and $\omega$ the frequency, known as the “boson peak” (BP). The microscopic origin of the BP remains one of the most controversial topics in condensed matter physics.
Some theories assume that BP arises from additional nonphononic excitations associated with quasi-localised modes (QLMs). The observation that structural glasses present nonphononic excitations is old [2]. This idea was inspired (see e.g. [3]) by the concept of resonant modes first formulated in the context of defective crystals [4] and it is a recurrent theme within the soft potential model of glassy anomalies [5, 6, 7, 8]. These nonphonic excitations have been widely reported in the low-frequency spectrum of amorphous solids and even disordered crystals [9], where they contribute to the density of states with a faster than Debye term, $D(\omega)\propto\omega^{p}$ with $p>(D-1)$, [10, 11, 12]. They have a quadrupolar nature and are characterized by a core made of few atoms accompanied by long-range displacements with power-law decay [13]. It has been argued that QLMs also exist at the boson peak frequency and cause the boson peak  [14, 7, 8, 15, 16, 17]. However, this scenario is difficult to prove, as QLMs would strongly hybridize with extended phononic modes around the BP frequency, rendering their identification [6, 18] and the study of their properties problematic. Henceforth, the existence of localised modes at the boson peak is questionable, motivating the development of alternative theories of the BP [19, 20, 21, 22, 23, 24, 25].
Recent works have explored the possibility of disentangling QLMs and extended modes at the boson peak frequency. In one approach, deformations with seemingly quadrupolar structures were studied within the vibrational modes at the boson peak frequency to assess whether they could serve as seeds for energy-minimal localized deformations [16, 17], which are identified with QLMs. Alternatively, Refs.[25, 26] examined how modes near the boson peak contribute to the amplitude of particle vibrational motion (see also[3]). Particles receiving significant contributions from these modes tend to form clusters, possibly emerging as a consequence of the hybridization of phonons and QLMs, and referred to as stringlets [25, 26] due to their resemblance to one-dimensional string-like or ring-like dynamical defects. Furthermore, the interaction between stringlets and phonons induces a flat boson peak mode [27, 28], observed in both simulations [25] and experiments [29]. This supports the hypothesis that stringlets, rather than quasi-localized quadrupolar modes, may be the microscopic origin of the boson peak, even in heated crystals [30]. It remains unclear whether localized defects are responsible for the boson peak in glasses and whether these defects are quasi-localized quadrupolar modes or string-like excitations, leaving the microscopic origin of the boson peak unresolved.
In this work, we resolve this tension by introducing a simple and efficient method based on cage-relative displacement to disentangle extended phonons from quasi-localized defects. We demonstrate that the modes at the boson peak frequency are a superposition of phonons and anisotropic, compact localized defects, exhibiting pure shear oscillatory deformations in two and three spatial dimensions. We show that the density of these defects explains the excess number of modes at the boson peak, providing strong support for previous suggestions that the boson peak originates from quadrupolar quasi-localized excitations [15, 16]. Additionally, we identify the so-called stringlets [25], revealing that, contrary to earlier speculations, they are not string-like objects but compact ones that align with the localized defects we have uncovered. Our work unifies different theoretical perspectives on the boson peak in glasses, resolving the long-standing debate and potentially establishing a consensus on its microscopic origin.
Figure 1: Attraction regulates the boson peak in 2D and 3D amorphous solids. a The reduced vDOS in three and (inset) two dimensions for systems differing in the extension $x_{c}$ of the intermolecular attractive potential well. In 3D, the materials lose stability as $x_{c}$ increases, as the boson peak intensity increases (b) while its frequency decreases (c). In 2D, attraction does not sensibly influence the excess vibrational modes. Frequencies are expressed in units of the natural frequency $\omega_{0}=c_{s}/a_{0}$, with $a_{0}$ the interparticle spacing and $c_{s}$ the shear wave speed. The inset of (c) illustrates the localised defects that, according to our investigation, induce the boson peak in a three-dimensional system with $x_{c}=1.7$.
Models and vibrational density of states
We numerically investigate the vibrational properties of two- and three-dimensional systems of polydisperse particles interacting via a Lennard-Jones-like potential $U(r_{ij},x_{c})$ which vanishes with its first two derivatives at $x_{c}\sigma$, $\sigma$ being the average particle diameter [31, 32]. We provide details in Methods. Previous works have shown that $x_{c}$, which controls the extension of the attractive well, influences the system’s relaxation dynamics [32] and the mechanical response [31, 33, 34, 35]. We simulate $N=256k$ particle systems in square or cubic simulation boxes with periodic boundary conditions at fixed interparticle spacing $a_{0}=\rho^{-1/d}\simeq 0.977$. We generate amorphous solid configurations by minimizing, via conjugate gradient, the energy of systems in thermal equilibrium in the NVT ensemble at $T=4.0\epsilon$, above the glass transition temperature for the considered $x_{c}$ values [32]. All data presented below are averaged over $200$ configurations at each $x_{c}$.
We evaluate the vibrational density of states (vDOS) over a broad frequency range by performing a Fourier transform of the velocity autocorrelation function in the linear response regime. Additionally, we compute the vDOS by directly diagonalizing the Hessian matrix, allowing us to access the eigenvectors, ${\bf e}_{k}$, and eigenfrequencies, $\omega_{k}$, in a low-frequency regime that includes the boson peak (BP) frequency. The two methods give consistent results (see SI). Fig. 1(a) illustrates the reduced-vDOS $D(\omega)/A_{D}\omega^{D}$, with $D_{\rm Debye}=A_{D}\omega^{D}$ Debye’s prediction (see SI), and demonstrates the BP in both two and three spatial dimensions. In 3D, the normalized peak amplitude increases with $x_{c}$ while its frequency $\omega_{\rm bp}$ decreases, indicating that shorter attractive wells lead to more stable glasses, as previously suggested [15, 33, 35, 31]. Due to the increase in stability limit we consider $x_{c}\geq 1.4\sigma$; at smaller $x_{c}$, phonons band structure persists to high frequency given our considered system sizes, making $\omega_{\rm bp}$ of difficult identification. In 2D, the role of $x_{c}$ is minor. In this case, we consider $x_{c}\geq 1.4\sigma$; at smaller $x_{c}$ the BP peak becomes very broad and $\omega_{\rm bp}$ of difficult estimation.
Disentangling phonons and localised defects
Due to the interaction with phonons, localised vibrations cannot be directly visualized as harmonic vibrations when their frequency is inside a phonon band. Instead, the phonons-defect hybridization leads to extended harmonic modes combining the two. Devising strategies to disentangle this hybridization and expose the localised defects is a longstanding challenge. Two approaches have been introduced so far: one relies on the assumption that the localised defects appear in regions where the local vibrational kinetic energy is high [3, 25, 26], and the other on the hypothesis that these defects have a quadrupolar structure visible in the hybridized field [16, 17].
We propose an alternative, simple and practical approach to disentangle localised defects and extended plane waves based on the hypothesis that plane waves at the BP frequency are locally affine while localised defects are not. Diverse approaches could be employed to identify regions of local non-affine behaviour [36]. Here we rely on the cage-relative displacement field, which has been previously used to filter out the effect of long-wavelength plane waves [37, 38, 39], to separate and remove the plane-wave contribution from the eigenvectors. This approach involves analyzing particles’ displacements relative to their immediate surroundings or the ‘cage’ formed by neighbouring particles. Specifically, given a displacement field ${\bm{u}}$, the cage-relative displacement ${\bm{u}}^{\leavevmode\nobreak\ \rm cr}_{i}$ of particle $i$ is ${\bm{u}}^{\leavevmode\nobreak\ \rm cr}_{i}={\bm{u}}_{i}-\frac{1}{N_{v}}\sum_{j=1}^{N_{v}}{\bm{u}}_{j}$, where the summation extends over the $N_{v}$ Voronoi neighbors of particle $i$. We use this approach by considering as displacement fields the vibrational modes of the systems, ${\bm{e}}_{k}$, which we determine via the direct diagonalization of the Hessian matrix. We consider as defective particles those with an $|{\bm{e}}^{\leavevmode\nobreak\ \rm cr}_{k,i}|^{2}$ value in the top 5%, unless otherwise specified, and determine the vibrational defects via a cluster analysis of these particles, considering two particles as belonging to the same cluster if Voronoi neighbors.
We validate our approach in Fig. 2. First, we demonstrate that this method does not filter out localized vibrations. In panel (a), we present a snapshot of a two-dimensional system where each particle is colour-coded based on its $|{\bm{e}}_{k,i}|^{2}$, for a low-frequency localised mode $k$. The localised excitation remains clearly visible in panel (b), where particles are coloured according to the magnitude of their cage-relative displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$, proving that our approach preserves localized defects. In panels (c) and (d), we compare the magnitude of the displacement and cage-relative displacement for the mode at the boson peak frequency. The cage-relative measure effectively filters out the collective translations responsible for the extended clusters in panel (c), revealing defective regions of large non-affine particle motion. This approach works equally well in three-dimensional systems, as shown in the inset of Fig. 1(c). Additional evidence in the Supplementary Information (SI) shows that cage-relative measures successfully filter out plane wave contributions at the boson peak frequency.
Figure 2: Disentangling extended plane waves and localised defects. a Colour map of the magnitude $|{\bm{e}}_{k,i}|^{2}$ of the particle displacement associated with a low-frequency mode with $\omega_{k}=0.23\omega_{\rm bp}$. A localized defect is visible. b The defect remains visible in the colour map of the magnitude of the cage-relative (CR) displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$. Panels c and d use the same approach to visualize the mode at the BP frequency. In this case, the cage-relative measure reveals the presence of many vibrational defects not apparent in the standard measure. In all plots, the magnitudes are scaled so that their average value is $1$.
Geometrical properties of vibrational defects
Fig. 3(a) shows that the probability distribution of the number of particles in a cluster $P(n)$ decays abruptly at small $n$ and roughly exponentially at larger $n$, similar to the distribution of the number of particles participating in the stringlets [40, 41, 25]. On increasing $x_{c}$, the average cluster size varies from $\langle n\rangle=16$ to $26$, in $3D$, and $\langle n\rangle=9$ to $13$, in $2D$. The average coordination number $z(n)$ grows with $n$ due to the competition between bulk and surface particles, as in Fig. 3(b), and it is larger than two if not for minute clusters, indicating that the clusters are not string-like. To assess a cluster’s geometrical properties, we perform a principal component analysis of the covariance matrix $C_{\alpha\beta}=\langle r_{\alpha}r_{\beta}\rangle-\langle r_{\alpha}\rangle\langle r_{\beta}\rangle$ of the positions of its particles. We use the eigenvalues $\lambda_{i}$ of $C_{\alpha\beta}$, which measure the variances of the particle positions along the principal directions, to define normalized standard deviations $\sigma_{i}=\sqrt{\lambda_{i}}/\sum_{i}\sqrt{\lambda_{i}}$ that obey the geometrical constraint $\sum_{i}\sigma_{i}=1$, and order them such that $\sigma_{i+1}>\sigma_{i}$. In two dimensions, $\sigma_{1}=1$ for linear (maximally anisotropic, one-dimensional) clusters and $\sigma_{1}=\frac{1}{2}$ for circular (isotropic) clusters. We find that the distribution $P(\sigma_{1})$ peaks at $\sigma_{1}\simeq 0.65$ and vanishes for $\sigma_{1}>0.8$, as shown in Fig. 3(c). This result further confirms that defects are not one-dimensional but rather asymmetric two-dimensional structures.
In three dimensions, for linear (1D), flat (2D), and spherical (isotropic 3D) clusters, $\sigma_{1}=1$, $\sigma_{1}+\sigma_{2}=1$, and $\sigma_{1}=\sigma_{2}=\sigma_{3}=\frac{1}{3}$, respectively. Fig. 3(d) presents density maps of the $P(\sigma_{1},\sigma_{2})$ probability distribution at $x_{c}=1.4$ and $x_{c}=1.7$ (shifted vertically) and suggests the localised defects have three-dimensional geometry with a variable degree of anisotropy.
Figure 3: Geometrical properties of vibrational defects. a Probability distribution of the number of particles belonging to each defect in 3D and 2D (inset). Symbols identify the values of $x_{c}$, as in Fig. 1b. b Average coordination number of the particles in clusters of size $n$. c We normalise ($\sum\sigma_{i}=1$) and order ($\sigma_{i}>\sigma_{i+1}$) the standard deviations of the particles’ positions along the principal axis of each ‘defect’, considering defects with at least $n=3$ particles. In two dimensions, the probability distribution of $P(\sigma_{1})$ peaks at intermediate values and drops at large ones, indicating that the defects are not one-dimensional ($\sigma_{1}=1$), but two-dimensional objects, albeit not precisely circular $(\sigma_{1}=1/2)$. d In three dimensions, the probability distribution $P(\sigma_{1},\sigma_{2})$ clarifies that clusters are not one-dimensional ($\sigma_{1}=1$) or two-dimensional $(\sigma_{1}+\sigma_{2}=1)$. We thus assume they are three-dimensional, albeit non-spherical $(\sigma_{1}=\sigma_{2}=1/3)$. Data for $x_{c}=1.7$ are shifted vertically for clarity.
Quadrupolar nature of vibrational defects
We obtain an insight into the vibrational nature of these defects by considering that the cage-relative displacement is conceptually related to a local strain tensor, as they are both insensitive to local uniform translations. We thus evaluate the local strain tensor ${\bm{\epsilon}}=\frac{1}{2}\left({\bm{\nabla}}{\bm{e}}+({\bm{\nabla}}\bm{e})^{T}\right),$ associated with the displacement field (eigenvector, ${\bm{e}}$) of the mode at the boson peak frequency, coarse-grained on grids of side-length $w\simeq 1.1$. We decompose this tensor into its volumetric and deviatoric components: $\bm{\varepsilon}=\bm{\varepsilon}_{\rm vol}+\bm{\varepsilon}_{\rm dev}$. $\bm{\varepsilon}_{\rm vol}$ represents local volumetric deformations. In contrast, the traceless deviatoric component represents volume-preserving deformations, which dominate the vibrational spectrum at the boson peak [42]. The eigenvalues $\upsilon_{k}$ of $\bm{\varepsilon}_{\rm dev}$ satisfy the traceless condition $\sum\upsilon_{k}=0$ and define a local quadrupolar charge $Q_{\rm loc}=\left(\frac{1}{2}\sum_{k=1}^{D}\upsilon_{k}^{2}\right)^{1/2}$ [43, 44].
Fig. 4(a) demonstrates a high Pearson correlation coefficient between the local quadrupolar charge and the coarse-grained cage-relative displacement magnitude, consistently with the expected analogy between local strain and cage-relative measure. Since defects are clusters of particles with significant coarse-grained cage-relative displacement, we calculate the deviatoric strain of a defect by summing the deviatoric strain of the individual particles within it. Specifically, we define a cluster deviatoric strain as $\mathcal{E}_{\rm dev}=\sum\bm{\varepsilon}_{\rm dev}$, where the sum extends over all bins containing particles of the considered cluster. The eigenvectors ${\bf U}_{k}$ and eigenvalues $\Upsilon_{k}$ of $\mathcal{E}_{\rm dev}$ describe the vibrational motion of the cluster. A cluster extends into the directions with $\Upsilon_{k}>0$ while it shrinks in the others for half period of oscillation, with particles reversing their direction of motion during the remaining half period. In the following, we assume $\Upsilon_{1}<0$ and order the eigenvalues so that $|\Upsilon_{i}|>|\Upsilon_{i+1}|$.
In two dimensions, the traceless condition implies $\Upsilon_{1}=-\Upsilon_{2}$. Figure 4(b) illustrates the quadrupoles associated with each cluster on a heat map of the magnitude of the coarse-grained cage-relative displacement, finding expected correlations in light of panel (a). The figure illustrates a small fraction of the system and represents each quadrupole with four arrows with a length proportional to ±$\pm\Upsilon_{i}{\bm{U}}_{i}$.
In three dimensions, we examine the relevance of various deformation modes that satisfy the traceless condition $\sum\Upsilon_{i}=0$ by analyzing the probability distribution $P(\Upsilon_{1},\Upsilon_{2})$. This distribution is supported within a range defined by the line $\Upsilon_{1}+\Upsilon_{2}=0$, which represents pure shear deformation, and the line $\Upsilon_{1}+2\Upsilon_{2}=0$, which corresponds to volume-preserving isotropic ($\Upsilon_{3}=\Upsilon_{2}=-\Upsilon_{1}/2$) uniaxial compressions. Figure 4(c) shows that quadrupolar deformations are predominantly pure shear, as $P(\Upsilon_{1},\Upsilon_{2})$ peaks near the $\Upsilon_{1}+\Upsilon_{2}=0$ line for different values of $x_{c}$. This result is further supported by direct visualization of the quadrupoles (Figure 4(d)), depicted as collections of 6 arrows with lengths proportional to ±$\pm\Upsilon_{i}{\bm{U}}_{i}$. Since $|\Upsilon_{3}|\ll|\Upsilon_{2}|$, the blue arrows corresponding to displacements in the ${\bm{U}_{3}}$ direction are not visible, making the quadrupoles appear as a two-dimensional cross.
Figure 4: Quadrupolar nature of defects. a Pearson correlation coefficient between coarse-grained cage-relative displacement and quadrupolar charge. b Quadrupoles associated with the cluster superimposed on a heat map of the magnitude of the coarse-grained cage-relative displacement, normalised by its average value. The figure refers to $x_{c}=1.7$ and depicts 1/12th of the system’s area. c Density map of the two largest (in modulus) eigenvalues of the deviatoric strain associated with the defects, normalised to its maximum value. The density is higher close to the $\Upsilon_{2}=-\Upsilon_{1}$ line, implying that the quadrupoles correspond to pure shear deformations. d We represent the quadrupolar deformations by attaching to each defect’s centre of mass the displacements along the principal axis ±$\pm\Upsilon_{1}{\bm{u}}_{1}$ (blue), ±$\pm\Upsilon_{2}{\bm{u}}_{2}$ (green) and ±$\pm\Upsilon_{3}{\bm{u}}_{3}$ (red). Since quadrupoles mostly correspond to pure-shear deformations, ${\bf u}_{3}$ is barely visible.
The localised defects regulate the boson peak
We test more rigorously whether quadrupolar defects regulate the boson peak by evaluating if their density $\rho_{d}$ (number per unit volume) accounts for the excess density of vibrational modes defined as $\Delta D(\omega_{\rm bp})=D(\omega_{\rm bp})-D_{\rm Debye}(\omega_{\rm bp})$, where $D_{\rm Debye}(\omega_{\rm bp})=A_{D}\omega^{D-1}$ represents Debye’s prediction, as expected if the phonon-defect interaction is weak. In making this test, it is important to note that $\rho_{d}$ is not uniquely determined, as it depends on the definition of defects used. We define defects as clusters of particles with cage-relative displacements in the top 5%. This threshold is not critical, as the SI shows that varying it does not significantly affect the number of defects or $\rho_{d}$, as it primarily impacts the defect size. Instead, the density of defects is more influenced by the minimum number of particles, $n_{\rm min}$, required for a cluster to be considered a defect. This threshold is introduced to ensure that volume-preserving deformations involve more than just a few particles.
Fig. 5(a) explores the dependence of the density of defects on $x_{c}$ in two-dimensional systems. An $n_{\rm min}$ increases, $\rho_{d}$ quickly looses any $x_{c}$ dependence. Similarly, we show in panel (b) that the excess of modes $\Delta D(\omega_{\rm bp})$ is also constant in two dimensions, showing therefore a clear correlation with the also constant $\rho_{d}$. In three dimensions, Fig. 5(c) demonstrates that the density of defects increases as $x_{c}$ decreases, consistent with the loss of stability. The excess of modes follows the same trend with $x_{c}$, with Fig. 5(d) demonstrating their proportionality. Importantly, in 2D, $\Delta D(\omega_{\rm bp})$ and $\rho_{d}$ have comparable values, besides being both $x_{c}$ independent. In 3D, we find that $\Delta D(\omega_{\rm bp})=c\rho_{d}$. The constant of proportionality $c$ depends on the specific definition of defects, but it is consistently of order 1. These findings strongly suggests that the excess of vibrational modes at the boson peak frequency originates from the quadrupolar defects we have identified.
Figure 5: Density of vibrational defects and excess density of modes. We evaluated the excess density of vibrational modes at the boson peak frequency, $\Delta D(\omega_{\rm bp})=D(\omega_{\rm bp})-D_{\rm Debye}(\omega_{\rm bp})$, $D_{\rm Debye}(\omega)$ being Debye’s density of state, and the density $\rho_{d}$ of vibrational defects containing more than $n$ particles. In two dimensions, a $\rho_{d}$ for clusters containing a minimum number of particle $n_{\rm min}$ is $x_{c}$ independent. Similarly, b, $\Delta D(\omega_{\rm bp})$ is $x_{c}$ independent. In three dimensions, c $\Delta D(\omega_{\rm bp})$ grows with $x_{c}$, and d $\rho_{d}\propto\Delta D(\omega_{\rm bp})$.
Discussion and relation to stringlets
We have identified localised defects hybridizing with plane waves at the boson peak frequency through a filtering technique based on the physical insight that localized vibrations are generally less affine than extended ones. This technique allowed us to isolate the vibrational defects and study their geometrical and vibrational properties. The vibrational defects have some degree of spatial anisotropy but are compact rather than planar or string-like. Their oscillatory motion is quadrupolar. Specifically, the defects undergo pure shear oscillatory deformations, extending along one axis while contracting along a perpendicular one in both two and three spatial dimensions. Analysis of various two- and three-dimensional systems shows that the number density of these defects quantitatively accounts for the excess of modes at the boson peak, confirming their physical relevance.
At this point, one might then wonder how our findings are compatible with the recent claims by Hu and Tanaka [25], that on the contrary identify the microscopic origin of the BP in one-dimensional string-like dynamical objects. These ‘stringlets‘ are defined as clusters of particles with a large vibrational amplitude, or equivalently, a high ‘vibrability’ $\Psi_{i}=\sum_{k}\frac{1}{\omega_{k}^{2}}\left|{\bf e}_{k,i}\right|^{2}$, where the sum extends over all modes with frequencies $\omega_{k}$ close to the boson peak frequency $|\omega_{\text{bp}}-\omega_{k}|<\Delta\omega$. We have performed the same stringlet analysis in our simulation model. As shown in Figure 6(a) by the average coordination number of the particles participating in the stringlets, we find that the clusters defined as in [25] are compact and extended objects, rather than one-dimensional, very similar to the localised defects found in our work (Figure 3(b)). In Figure 6(b), we observe a high spatial correlation between particles with vibrability in the top 5% (red) alongside those with cage-relative displacements in the top 5% (blue), indicating a high correlation between vibrational motion and cage-relative displacement at the BP frequency. These correlations are further confirmed by measuring the fraction $f$ of particles with high vibrability and cage-relative displacement, which would be $f_{\text{rand}}=0.05$ in the absence of correlations. Figure 6(c) shows that $f$ is independent of $x_{c}$ and significantly exceeds the random expectation.
Our analysis provides strong evidence that quadrupolar localised defects and high-vibrability clusters (previously identified as stringlets) are two sides of the same coin. This reconciles two of the most successful theories for the origin of the BP in glasses and reveal the ultimate nature of the defects responsible for it.
Along the lines of [15], it would be interesting to understand whether the same defective structures could be also rationalized by means of shear modulus fluctuations and whether our findings are compatible with another successful theory of the BP – heterogeneous elasticity theory [45]. Moreover, it would be fruitful to investigate the relation between the defects discussed in our manuscript and (I) the geometric charges responsible for plastic screening and anomalous elastic response [46], (II) the topological vortex-like defects recently connected to plastic soft spots [47, 48].
Figure 6: Reconciling stringlets and quadrupolar defects. a The coordination number of vibrability clusters increases with $n$, following a trend similar to that of clusters formed based on cage-relative displacement (Fig. 3c). b Particles exhibiting high cage-relative displacement at the boson peak (red) and high vibrability (blue), both in the top 5%, for $x_{c}=1.7$. Strong correlations between the two are observed. c The fraction of particles displaying both large cage-relative displacement and high vibrability significantly exceeds the expectation under random conditions.
Methods
Numerical model
We consider two- and three-dimensional systems of particles of diameter $\sigma_{i}$, drawn from a uniform random distribution in the range [0.8:1.2], that interact via family of LJ-like potentials consisting of a repulsive and an attractive part, $U(r_{ij},x_{c})=U_{r}(r_{ij})+U_{a}(r_{ij},x_{c})$ [31, 32]. The repulsive part follows the standard LJ functional form,
$U_{r}(r_{ij})=4\epsilon\left[\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{12}-\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}\right],$ (1)
where $\sigma_{ij}=(\sigma_{i}+\sigma_{j})/2$, but only acts for $r_{ij}\leq r_{ij}^{\rm min}=2^{1/6}\sigma_{ij}$. The attractive part, which only acts for distances in the range $2^{1/6}\sigma_{ij}\leq r_{ij}\leq x_{c}\sigma_{ij}$, is given by
$U_{a}(r_{ij})=\epsilon\left[a_{0}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{12}-a_{1}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}+\sum_{l=0}^{3}{c_{2l}\left(\frac{r_{ij}}{\sigma_{ij}}\right)^{2l}}\right].$ (2)
The parameters $a_{0},a_{1}$ and $c_{2l}$ are chosen such that the potential $U(r_{ij})$ and its first two derivatives are continuous at the minimum $r_{ij}^{\rm min}$ and at the cutoff $r_{ij}^{(c)}=x_{c}\sigma_{ij}$, where the potential also vanishes. The parameter $x_{c}$ sets the extension of the attractive well [31], which vanishes at $x_{c}\sigma$. This parameter influences the relaxation dynamics [32] and the mechanical response [31, 33, 34, 35].
Data availability
The datasets generated and analysed during the current study are available upon reasonable request by contacting the corresponding authors.
Funding
MPC acknowledges support by the Singapore Ministry of Education through grants MOE-T2EP50221-0016 and RG152/23. CJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). MB acknowledges the support of the sponsorship from the Yangyang Development Fund.
Author contributions statement
M.P.C. and M.B. devised the project. M.P.C., S.M. and M.B. wrote the paper. S.M., D.S.Y.H. performed the numerical simulations and prepared the figures. All authors contributed to the scientific discussion and refinement of the manuscript.
Supplementary information
Density of vibrational modes
We compare the density of vibrational modes obtained from Fourier transforming the velocity autocorrelation function (points) and via the direct diagonalisation of the Hessian matrix in Fig. 7 for systems with 256k particles. In the figure, we also illustrate the Debye’s prediction, $D_{\rm Debye}=A_{D}\omega^{D-1}$. In 2D, $A_{D}=\frac{2}{8\pi\rho/(c_{l}^{-2}+c_{s}^{-2})}$, with $\rho$ the number density and $c_{l}=\sqrt{\frac{K+\mu}{\rho}}$ and $c_{s}=\sqrt{\frac{\mu}{\rho}}$ the longitudinal and the transverse shear velocities. In 3D, $A_{D}=\frac{3}{18\pi^{2}\rho/(c_{l}^{-3}+2c_{s}^{-3})}$, $c_{l}=\sqrt{\frac{K+\frac{4\mu}{3}}{\rho}}$ and $c_{s}=\sqrt{\frac{\mu}{\rho}}$. We determine the bulk $K$ and shear $\mu$ moduli by investigating the stress-strain relationships in the linear response regime.
Figure 7: Different approaches to evaluate the vDOS. We compare diverse approaches to evaluating the vDOS, in 2D and 3D. Evaluating the vDOS by Fourier transforming the velocity autocorrelation function and via the diagonalization of the Hessian gives consistent results. Due to our system’s large size, 256k particles, we compute only the lower end of the spectrum via the direct diagonalization of the Hessian.
Cage-relative measures filter plane waves at the boson peak
To evaluate the cage-relative ability to filter extended modes of norm $|{\bm{u}}|^{2}=1$, we consider $|{\bm{u}}^{\rm cr}|^{2}$. Filtering occurs if $|{\bm{u}}^{\rm cr}|^{2}\ll 1$. As extended modes, we consider phonons with various wavelengths. There is no need to consider the superposition of phonons as the cage-relative ‘operator’ is linear, $(a{\bm{u}}+b{\bm{v}})^{\rm cr}=a{\bm{u}}^{\rm cr}+b{\bm{v}}^{\rm cr}$. Any linear combination of ‘filtered’ displacement fields is also filtered.
Fig. 8 illustrates $|{\bm{p}}^{\rm cr}|^{2}$ for phonon ${\bf p}$ of wavelength $\lambda$, in two dimensions. The larger the phonon wavelength, the more efficient the filtering. Importantly, filtering is highly efficient at the wavelength $\lambda_{\rm bp}=\frac{2\pi c_{s}}{\omega_{\rm bp}}$ of the phonons expected to hybridize at the boson peak, as $|{\bm{p}}^{\rm cr}|^{2}$ is in the range $10^{-4}$ to $10^{-3}$, its value depending on the cutoff $x_{c}$ of the interaction potential.
Figure 8: Filtering phonons via the cage-relative measured. Squared norm of the cage-relative displacement field $|{\bm{p}}^{\rm cr}|^{2}$ associated with phonons ($|{\bm{p}}|^{2}=1$) of wavelength $\lambda$, in two dimensions. The vertical lines mark the wavelength $\lambda_{\rm bp}$ of phonons at the boson peak frequency for the extremal investigated $x_{c}$ values. Analogous results hold in 3D.
Density of vibrational defects
We define defects as clusters of particles whose cage-relative displacement is within the top $c=5\%$. We explore how the density of defects, $\rho_{d}$, depends on the percentage $c$ of particles with the highest cage-relative displacement.
As $c\to 0$, $\rho_{d}$ vanishes since there are no clusters, meaning no particles meet the clustering criteria. As $c$ increases, $\rho_{d}$ grows, reflecting both the increasing size and number of clusters. However, beyond a certain point, as clusters begin to merge, $\rho_{d}$ decreases, leading to a maximum in defect density at an intermediate $c$ value.
Fig. 6a shows that, in 2D, the density of defects remains $x_{c}$ independent for $c<10\%$. Unless this threshold is considered unrealistically high, the exact value of $c$ does not significantly affect our conclusions.
In 3D, the behavior of $\rho_{d}$ for small $x_{c}$ values is illustrated in Fig. 6c: $\rho_{d}$ remains unchanged between $x_{c}=1.8$ and $x_{c}=1.7$ and increases as $x_{c}$ decreases further. Thus, our results are not highly sensitive to the exact choice of $x_{c}$.
Figure 9: Cutoff dependence of the density of defects. We perform a cluster analysis of the cof the particles with the largest magnitude of the cage relative displacement (in the mode at boson peak frequency) to identify the defects. a the density of defects $\rho_{d}$ as a maximum as a function of $c$, and is roughly constant as $c$ varies from $5\%$ to $7\%$. b $\rho_{d}$ remains constant in this $c$ range for all $x_{c}$, also when we consider as defects only cluster with at least $n_{\rm min}$ particles. We illustrate the considered values of $n_{\rm min}$ for $x_{c}=1.5$.

### v2

@@
-Revealing the Geometrical and Vibrational Properties of the Defects Driving the Boson Peak
+Geometrical and Vibrational Properties of the Defects Driving the Boson Peak
 Abstract
-The vibrational density of states is key to understanding the mechanical, thermal, and transport properties of materials.
-In amorphous solids, this density shows an excess of vibrational modes compared to the Debye model, known as the boson peak, whose origin remains poorly understood.
-Previous studies have suggested a link between the boson peak and quasi-localized nonphononic vibrations, or "defects." However, it has been difficult to clearly identify these defects, possibly because they hybridize with extended phonons, casting doubt on their existence and connection to the boson peak.
-In this work, we introduce a simple and practical method for separating hybridized phonons from localized vibrations.
-We show that phonons at the boson peak frequency hybridize with localized defects.
-These defects are anisotropic, compact, and exhibit oscillatory pure shear deformations.
-Their density correlates with the excess of vibrational modes at the boson peak frequency across various two- and three-dimensional systems, confirming that they are the microscopic origin of the boson peak.
+In amorphous solids, the vibrational density of states shows an excess of modes over the Debye model, known as the boson peak, whose origin remains unclear.
+Studies suggest a link to quasi-localized nonphononic vibrations or ‘defects,’ but identifying them is challenging due to hybridization with phonons that renders methods based on localization properties, such as the participation ratio, unreliable.
+We introduce a practical method to separate hybridized phonons from localized vibrations and find that boson peak phonons hybridize with compact, two-dimensional defects exhibiting oscillatory pure shear deformations.
+These two-dimensional defects are also exposed by the procedure recently employed to identify stringlets (Nature Physics volume 18, pages 669–677 (2022)), suggesting that these may not be one-dimensional objects as speculated.
+Our work demonstrates the presence of localized defects at the boson peak frequency and provides a comprehensive characterization of their vibrational and geometric properties, resolving the tension between the concepts of quasi-localized quadrupolar defects and stringlets.
 Thermal and vibrational properties of amorphous solids display universal characteristics that are very different from their crystalline counterparts [1].
-In particular, their vibrational density of states (vDOS) $D(\omega)$ presents a universal excess anomaly over the Debye prediction $D(\omega)\propto\omega^{D-1}$, with $D$ the number of spatial dimensions and $\omega$ the frequency, known as the “boson peak” (BP).
+In particular, their vibrational density of states (vDOS) $D(\omega)$ presents a universal excess anomaly over the Debye prediction $\color[rgb]{0,0,0}D(\omega)\propto\omega^{d-1}$}, with $\color[rgb]{0,0,0}d$} the number of spatial dimensions and $\omega$ the frequency, known as the “boson peak” (BP).
 The microscopic origin of the BP remains one of the most controversial topics in condensed matter physics.
@@
 This idea was inspired (see e.g.
-[3]) by the concept of resonant modes first formulated in the context of defective crystals [4] and it is a recurrent theme within the soft potential model of glassy anomalies [5, 6, 7, 8].
-These nonphonic excitations have been widely reported in the low-frequency spectrum of amorphous solids and even disordered crystals [9], where they contribute to the density of states with a faster than Debye term, $D(\omega)\propto\omega^{p}$ with $p>(D-1)$, [10, 11, 12].
+[3, 4]) by the concept of resonant modes first formulated in the context of defective crystals [5] and it is a recurrent theme within the soft potential model of glassy anomalies [6, 7, 8, 9].
+These nonphononic excitations have been widely reported in the low-frequency spectrum of amorphous solids and even disordered crystals [10], where they contribute to the density of states with a faster than Debye term, $D(\omega)\propto\omega^{p}$ with $\color[rgb]{0,0,0}p>(d-1)$}[11, 12, 13, 14, 15].
 They have a quadrupolar nature and are characterized by a core made of few atoms accompanied by long-range displacements with power-law decay [16].
-It has been argued that QLMs also exist at the boson peak frequency and cause the boson peak  [14, 7, 8, 15, 16, 17].
-However, this scenario is difficult to prove, as QLMs would strongly hybridize with extended phononic modes around the BP frequency, rendering their identification [6, 18] and the study of their properties problematic.
-Henceforth, the existence of localised modes at the boson peak is questionable, motivating the development of alternative theories of the BP [19, 20, 21, 22, 23, 24, 25].
-Recent works have explored the possibility of disentangling QLMs and extended modes at the boson peak frequency.
-In one approach, deformations with seemingly quadrupolar structures were studied within the vibrational modes at the boson peak frequency to assess whether they could serve as seeds for energy-minimal localized deformations [16, 17], which are identified with QLMs.
-Alternatively, Refs.[25, 26] examined how modes near the boson peak contribute to the amplitude of particle vibrational motion (see also[3]).
-Particles receiving significant contributions from these modes tend to form clusters, possibly emerging as a consequence of the hybridization of phonons and QLMs, and referred to as stringlets [25, 26] due to their resemblance to one-dimensional string-like or ring-like dynamical defects.
-Furthermore, the interaction between stringlets and phonons induces a flat boson peak mode [27, 28], observed in both simulations [25] and experiments [29].
-This supports the hypothesis that stringlets, rather than quasi-localized quadrupolar modes, may be the microscopic origin of the boson peak, even in heated crystals [30].
-It remains unclear whether localized defects are responsible for the boson peak in glasses and whether these defects are quasi-localized quadrupolar modes or string-like excitations, leaving the microscopic origin of the boson peak unresolved.
-In this work, we resolve this tension by introducing a simple and efficient method based on cage-relative displacement to disentangle extended phonons from quasi-localized defects.
-We demonstrate that the modes at the boson peak frequency are a superposition of phonons and anisotropic, compact localized defects, exhibiting pure shear oscillatory deformations in two and three spatial dimensions.
-We show that the density of these defects explains the excess number of modes at the boson peak, providing strong support for previous suggestions that the boson peak originates from quadrupolar quasi-localized excitations [15, 16].
+It has been argued that similar defects also exist at higher frequencies and cause the boson peak [17, 8, 9, 18, 19, 20].
+However, this scenario is difficult to prove, as these defects would strongly hybridize with extended phononic modes around the BP frequency.
+In particular, methods based on localization properties, such as the participation ratio, are not effective at these frequencies.
+If defect exists at high frequencies, they do not appear as QLM, rendering their identification [7, 21] and the study of their properties problematic.
+Henceforth, the existence of localised defects at the boson peak is questionable, motivating the development of alternative theories of the BP [22, 23, 24, 25, 26, 27, 28].
+Recent works have explored the possibility of disentangling localised defects and extended modes at the boson peak frequency.
+In one approach, deformations with seemingly quadrupolar structures were visually isolated from vibrational modes at the boson peak frequency to assess whether they could serve as seeds for energy-minimal localized deformations [19, 20], which are identified with QLMs.
+However, local quadrupolar deformations are not generally apparent in the hybridized displacement field, as demonstrated in Appendix A, and defects might, in principle, induce diverse localized deformations.
+Alternatively, one could examine how modes near the boson peak contribute to the amplitude of particle vibrational motion [3].
+Refs.
+[28, 29] have suggested that particles receiving significant contributions from these modes tend to organize along string-like clusters or stringlets with one-dimensional nature and lacking quadrupolar characteristics.
+The interaction between stringlets and phonons induces a flat boson peak mode [30, 31], observed in both simulations [28] and experiments [32] supports the hypothesis that stringlets, rather than quasi-localized quadrupolar modes, could be the microscopic origin of the boson peak, even in heated crystals [33].
+It remains unclear whether localized defects are responsible for the boson peak in glasses and whether these defects are quasi-localized quadrupolar modes or string-like excitations.
+Thus, the microscopic origin of the boson peak remains unresolved.
+In this work, we resolve this tension by introducing a simple and efficient method based on cage-relative (CR) displacement to disentangle extended phonons from quasi-localized defects.
+We demonstrate that the modes at the boson peak frequency are a superposition of phonons and anisotropic, compact, two-dimensional localized defects, exhibiting pure shear oscillatory deformations in two and three spatial dimensions.
+We show that the size of these defects correlates with the length scales associated with the boson peak, providing strong support for previous suggestions that the boson peak originates from quadrupolar quasi-localized excitations [18, 19].
 Additionally, we identify the so-called stringlets [28], revealing that, contrary to earlier speculations, they are not string-like objects but compact ones that align with the localized defects we have uncovered.
 Our work unifies different theoretical perspectives on the boson peak in glasses, resolving the long-standing debate and potentially establishing a consensus on its microscopic origin.
-Figure 1: Attraction regulates the boson peak in 2D and 3D amorphous solids.
-a The reduced vDOS in three and (inset) two dimensions for systems differing in the extension $x_{c}$ of the intermolecular attractive potential well.
-In 3D, the materials lose stability as $x_{c}$ increases, as the boson peak intensity increases (b) while its frequency decreases (c).
-In 2D, attraction does not sensibly influence the excess vibrational modes.
-Frequencies are expressed in units of the natural frequency $\omega_{0}=c_{s}/a_{0}$, with $a_{0}$ the interparticle spacing and $c_{s}$ the shear wave speed.
-The inset of (c) illustrates the localised defects that, according to our investigation, induce the boson peak in a three-dimensional system with $x_{c}=1.7$.
-Models and vibrational density of states
-We numerically investigate the vibrational properties of two- and three-dimensional systems of polydisperse particles interacting via a Lennard-Jones-like potential $U(r_{ij},x_{c})$ which vanishes with its first two derivatives at $x_{c}\sigma$, $\sigma$ being the average particle diameter [31, 32].
-We provide details in Methods.
+We numerically investigate the vibrational properties of two- and three-dimensional systems of polydisperse particles interacting via a Lennard-Jones-like potential $U(r_{ij},x_{c})$ which vanishes with its first two derivatives at $x_{c}\sigma$, $\sigma$ being the average particle diameter of the interacting particles [34, 35].
 Previous works have shown that $x_{c}$, which controls the extension of the attractive well, influences the system’s relaxation dynamics [35] and the mechanical response [34, 36, 37, 38].
-We simulate $N=256k$ particle systems in square or cubic simulation boxes with periodic boundary conditions at fixed interparticle spacing $a_{0}=\rho^{-1/d}\simeq 0.977$.
+We simulate systems with $N=32k$ particles - which are enough for finite size systems to be negligible at the boson peak frequency [39, 29, 19] - in square or cubic simulation boxes with periodic boundary conditions at fixed interparticle spacing $a_{0}=\rho^{-1/D}\simeq 0.977$.
 We generate amorphous solid configurations by minimizing, via conjugate gradient, the energy of systems in thermal equilibrium in the NVT ensemble at $T=4.0\epsilon$, above the glass transition temperature for the considered $x_{c}$ values [35].
 All data presented below are averaged over $200$ configurations at each $x_{c}$.
-We evaluate the vibrational density of states (vDOS) over a broad frequency range by performing a Fourier transform of the velocity autocorrelation function in the linear response regime.
-Additionally, we compute the vDOS by directly diagonalizing the Hessian matrix, allowing us to access the eigenvectors, ${\bf e}_{k}$, and eigenfrequencies, $\omega_{k}$, in a low-frequency regime that includes the boson peak (BP) frequency.
-The two methods give consistent results (see SI).
-Fig.
-1(a) illustrates the reduced-vDOS $D(\omega)/A_{D}\omega^{D}$, with $D_{\rm Debye}=A_{D}\omega^{D}$ Debye’s prediction (see SI), and demonstrates the BP in both two and three spatial dimensions.
-In 3D, the normalized peak amplitude increases with $x_{c}$ while its frequency $\omega_{\rm bp}$ decreases, indicating that shorter attractive wells lead to more stable glasses, as previously suggested [15, 33, 35, 31].
-Due to the increase in stability limit we consider $x_{c}\geq 1.4\sigma$; at smaller $x_{c}$, phonons band structure persists to high frequency given our considered system sizes, making $\omega_{\rm bp}$ of difficult identification.
-In 2D, the role of $x_{c}$ is minor.
-In this case, we consider $x_{c}\geq 1.4\sigma$; at smaller $x_{c}$ the BP peak becomes very broad and $\omega_{\rm bp}$ of difficult estimation.
-Disentangling phonons and localised defects
-Due to the interaction with phonons, localised vibrations cannot be directly visualized as harmonic vibrations when their frequency is inside a phonon band.
-Instead, the phonons-defect hybridization leads to extended harmonic modes combining the two.
-Devising strategies to disentangle this hybridization and expose the localised defects is a longstanding challenge.
-Two approaches have been introduced so far: one relies on the assumption that the localised defects appear in regions where the local vibrational kinetic energy is high [3, 25, 26], and the other on the hypothesis that these defects have a quadrupolar structure visible in the hybridized field [16, 17].
-We propose an alternative, simple and practical approach to disentangle localised defects and extended plane waves based on the hypothesis that plane waves at the BP frequency are locally affine while localised defects are not.
-Diverse approaches could be employed to identify regions of local non-affine behaviour [36].
-Here we rely on the cage-relative displacement field, which has been previously used to filter out the effect of long-wavelength plane waves [37, 38, 39], to separate and remove the plane-wave contribution from the eigenvectors.
+We provide details on the model, the numerical procedures, and the $x_{c}$ dependence of the vibrational model in Appendix B.
+Exposing hybridized defects – We propose a simple and practical approach to disentangle localised defects and extended plane waves based on the hypothesis that plane waves at the BP frequency are locally affine while localised defects are not.
+Diverse approaches could be employed to decompose a vibration into a phononic and a non-phononic component [40, 41], or more generally, to identify regions of local non-affine behaviour [42, 43].
+Here, we rely on the cage-relative (CR) displacement field to separate and remove the plane-wave contribution from the eigenvectors.
+This approach has been previously used to filter out the effect of low-frequency long-wavelength plane waves on the low-temperature liquid dynamics [44, 45, 46], never to isolate defects.
 This approach involves analyzing particles’ displacements relative to their immediate surroundings or the ‘cage’ formed by neighbouring particles.
-Specifically, given a displacement field ${\bm{u}}$, the cage-relative displacement ${\bm{u}}^{\leavevmode\nobreak\ \rm cr}_{i}$ of particle $i$ is ${\bm{u}}^{\leavevmode\nobreak\ \rm cr}_{i}={\bm{u}}_{i}-\frac{1}{N_{v}}\sum_{j=1}^{N_{v}}{\bm{u}}_{j}$, where the summation extends over the $N_{v}$ Voronoi neighbors of particle $i$.
-We use this approach by considering as displacement fields the vibrational modes of the systems, ${\bm{e}}_{k}$, which we determine via the direct diagonalization of the Hessian matrix.
-We consider as defective particles those with an $|{\bm{e}}^{\leavevmode\nobreak\ \rm cr}_{k,i}|^{2}$ value in the top 5%, unless otherwise specified, and determine the vibrational defects via a cluster analysis of these particles, considering two particles as belonging to the same cluster if Voronoi neighbors.
-We validate our approach in Fig.
-2.
-First, we demonstrate that this method does not filter out localized vibrations.
-In panel (a), we present a snapshot of a two-dimensional system where each particle is colour-coded based on its $|{\bm{e}}_{k,i}|^{2}$, for a low-frequency localised mode $k$.
-The localised excitation remains clearly visible in panel (b), where particles are coloured according to the magnitude of their cage-relative displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$, proving that our approach preserves localized defects.
-In panels (c) and (d), we compare the magnitude of the displacement and cage-relative displacement for the mode at the boson peak frequency.
-The cage-relative measure effectively filters out the collective translations responsible for the extended clusters in panel (c), revealing defective regions of large non-affine particle motion.
+Specifically, given a displacement field $\bm{u}$}, the cage-relative displacement ${\bm{u}}^{\rm cr}_{i}$ of particle $i$ is ${\bm{u}}^{\rm cr}_{i}={\bm{u}}_{i}-\frac{1}{N_{v}}\sum_{j=1}^{N_{v}}{\bm{u}}_{j}$, where the summation extends over the $N_{v}$ voronoi neighbors of particle $i$.
+We use this approach by considering as displacement fields the vibrational modes of the systems, ${\bm{e}}_{k}$, which we determine through the direct diagonalization of the Hessian matrix.
+We validate this approach in Appendix C by demonstrating that the CR measure preserves the localized defects and successfully filters out the plane wave contributions at the boson peak frequency.
+As an example, Fig.
+1 illustrates a snapshot of a two-dimensional system where each particle is colour-coded based on the magnitude of the displacement $|{\bm{e}}_{k,i}|^{2}$ (in panel a), and cage-relative displacement $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$ (in panel b), for an eigenmode at the boson peak frequency.
+The cage-relative measure reveals defective regions of large non-affine particle motion.
 This approach works equally well in three-dimensional systems, as shown in the inset of Fig.
-1(c).
-Additional evidence in the Supplementary Information (SI) shows that cage-relative measures successfully filter out plane wave contributions at the boson peak frequency.
-Figure 2: Disentangling extended plane waves and localised defects.
-a Colour map of the magnitude $|{\bm{e}}_{k,i}|^{2}$ of the particle displacement associated with a low-frequency mode with $\omega_{k}=0.23\omega_{\rm bp}$.
-A localized defect is visible.
-b The defect remains visible in the colour map of the magnitude of the cage-relative (CR) displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$.
-Panels c and d use the same approach to visualize the mode at the BP frequency.
-In this case, the cage-relative measure reveals the presence of many vibrational defects not apparent in the standard measure.
-In all plots, the magnitudes are scaled so that their average value is $1$.
-Geometrical properties of vibrational defects
-Fig.
-3(a) shows that the probability distribution of the number of particles in a cluster $P(n)$ decays abruptly at small $n$ and roughly exponentially at larger $n$, similar to the distribution of the number of particles participating in the stringlets [40, 41, 25].
-On increasing $x_{c}$, the average cluster size varies from $\langle n\rangle=16$ to $26$, in $3D$, and $\langle n\rangle=9$ to $13$, in $2D$.
-The average coordination number $z(n)$ grows with $n$ due to the competition between bulk and surface particles, as in Fig.
-3(b), and it is larger than two if not for minute clusters, indicating that the clusters are not string-like.
-To assess a cluster’s geometrical properties, we perform a principal component analysis of the covariance matrix $C_{\alpha\beta}=\langle r_{\alpha}r_{\beta}\rangle-\langle r_{\alpha}\rangle\langle r_{\beta}\rangle$ of the positions of its particles.
-We use the eigenvalues $\lambda_{i}$ of $C_{\alpha\beta}$, which measure the variances of the particle positions along the principal directions, to define normalized standard deviations $\sigma_{i}=\sqrt{\lambda_{i}}/\sum_{i}\sqrt{\lambda_{i}}$ that obey the geometrical constraint $\sum_{i}\sigma_{i}=1$, and order them such that $\sigma_{i+1}>\sigma_{i}$.
-In two dimensions, $\sigma_{1}=1$ for linear (maximally anisotropic, one-dimensional) clusters and $\sigma_{1}=\frac{1}{2}$ for circular (isotropic) clusters.
-We find that the distribution $P(\sigma_{1})$ peaks at $\sigma_{1}\simeq 0.65$ and vanishes for $\sigma_{1}>0.8$, as shown in Fig.
-3(c).
-This result further confirms that defects are not one-dimensional but rather asymmetric two-dimensional structures.
-In three dimensions, for linear (1D), flat (2D), and spherical (isotropic 3D) clusters, $\sigma_{1}=1$, $\sigma_{1}+\sigma_{2}=1$, and $\sigma_{1}=\sigma_{2}=\sigma_{3}=\frac{1}{3}$, respectively.
-Fig.
-3(d) presents density maps of the $P(\sigma_{1},\sigma_{2})$ probability distribution at $x_{c}=1.4$ and $x_{c}=1.7$ (shifted vertically) and suggests the localised defects have three-dimensional geometry with a variable degree of anisotropy.
-Figure 3: Geometrical properties of vibrational defects.
-a Probability distribution of the number of particles belonging to each defect in 3D and 2D (inset).
-Symbols identify the values of $x_{c}$, as in Fig.
-1b.
+6(c).
+We consider as defective particles those with an $|{\bm{e}}^{\penalty 10000\ \rm cr}_{k,i}|^{2}$ value in the top 5%, unless otherwise specified, and determine the vibrational defects via a cluster analysis of these particles, considering two particles as belonging to the same cluster if voronoi neighbours.
+In the following, we discuss results obtained via the analysis of a single mode at the BP frequency, for each system.
+We have checked that analogous results occur for close frequencies $|\omega-\omega_{\rm bp}|<0.1\omega_{\rm bp}$, but not at higher ones.
+Figure 1: a Colour map of the magnitude $|{\bm{e}}_{k,i}|^{2}$ of the particle displacement associated with an eigenmode with $\omega_{k}=\omega_{\rm bp}$.
+Swirls characterizing hybridization with plane waves are observed.
+b The cage-relative displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$ reveals the presence of many vibrational defects not apparent in the standard measure.
+In these plots, the magnitudes are scaled so that their average value is 11.
+Figure 2: a Probability distribution of the number of particles belonging to each defect in 3D and 2D (inset).
 b Average coordination number of the particles in clusters of size $n$.
@@
 Data for $x_{c}=1.7$ are shifted vertically for clarity.
-Quadrupolar nature of vibrational defects
-We obtain an insight into the vibrational nature of these defects by considering that the cage-relative displacement is conceptually related to a local strain tensor, as they are both insensitive to local uniform translations.
-We thus evaluate the local strain tensor ${\bm{\epsilon}}=\frac{1}{2}\left({\bm{\nabla}}{\bm{e}}+({\bm{\nabla}}\bm{e})^{T}\right),$ associated with the displacement field (eigenvector, ${\bm{e}}$) of the mode at the boson peak frequency, coarse-grained on grids of side-length $w\simeq 1.1$.
-We decompose this tensor into its volumetric and deviatoric components: $\bm{\varepsilon}=\bm{\varepsilon}_{\rm vol}+\bm{\varepsilon}_{\rm dev}$.
-$\bm{\varepsilon}_{\rm vol}$ represents local volumetric deformations.
-In contrast, the traceless deviatoric component represents volume-preserving deformations, which dominate the vibrational spectrum at the boson peak [42].
-The eigenvalues $\upsilon_{k}$ of $\bm{\varepsilon}_{\rm dev}$ satisfy the traceless condition $\sum\upsilon_{k}=0$ and define a local quadrupolar charge $Q_{\rm loc}=\left(\frac{1}{2}\sum_{k=1}^{D}\upsilon_{k}^{2}\right)^{1/2}$ [43, 44].
+e Radius of gyration, $R_{g}$, as a function of cluster size $n$.
+3D data are shifted by a factor 33 for clarity.
+f The characteristic length scale $\lambda$ of defects governs the boson peak frequency in two and three dimensions.
+In all panels, symbols identify the values of $x_{c}$.
+Geometrical properties – Fig.
+2(a) shows that the probability distribution of the number of particles in a cluster $P(n)$ decays abruptly at small $n$ and roughly exponentially at larger $n$, similar to the distribution of the number of particles participating in the stringlets [47, 48, 28].
+On increasing $x_{c}$, the average cluster size varies from $\langle n\rangle=16$ to $26$, in 3D, and $\langle n\rangle=9$ to $13$, in 2D.
+The average coordination number $\color[rgb]{0,0,0}Z(n)$} grows with $n$ due to the competition between bulk and surface particles, as in Fig.
+2(b), and it is larger than two if not for minute clusters, indicating that the clusters are not string-like.
+To further assess a cluster’s geometrical properties, we perform a principal component analysis of the covariance matrix $C_{\alpha\beta}=\langle r_{\alpha}r_{\beta}\rangle-\langle r_{\alpha}\rangle\langle r_{\beta}\rangle$ of the positions of its particles.
+We use the eigenvalues $\lambda_{i}$ of $C_{\alpha\beta}$, which measure the variances of the particle positions along the principal directions, to define normalized standard deviations $\sigma_{i}=\sqrt{\lambda_{i}}/\sum_{i}\sqrt{\lambda_{i}}$ that obey the geometrical constraint $\sum_{i}\sigma_{i}=1$, and order them such that $\sigma_{i}>\sigma_{i+1}$.
+In two dimensions, $\sigma_{1}=1$ for linear (maximally anisotropic, one-dimensional) clusters and $\sigma_{1}=\frac{1}{2}$ for circular (isotropic) clusters.
+We find that the distribution $P(\sigma_{1})$ peaks at $\sigma_{1}\simeq 0.65$ and vanishes for $\sigma_{1}>0.8$, as shown in Fig.
+2(c).
+In three dimensions, for linear (1D), flat (2D), and spherical (isotropic 3D) clusters, $\sigma_{1}=1$, $\sigma_{1}+\sigma_{2}=1$, and $\sigma_{1}=\sigma_{2}=\sigma_{3}=\frac{1}{3}$, respectively.
 Fig.
-4(a) demonstrates a high Pearson correlation coefficient between the local quadrupolar charge and the coarse-grained cage-relative displacement magnitude, consistently with the expected analogy between local strain and cage-relative measure.
-Since defects are clusters of particles with significant coarse-grained cage-relative displacement, we calculate the deviatoric strain of a defect by summing the deviatoric strain of the individual particles within it.
-Specifically, we define a cluster deviatoric strain as $\mathcal{E}_{\rm dev}=\sum\bm{\varepsilon}_{\rm dev}$, where the sum extends over all bins containing particles of the considered cluster.
-The eigenvectors ${\bf U}_{k}$ and eigenvalues $\Upsilon_{k}$ of $\mathcal{E}_{\rm dev}$ describe the vibrational motion of the cluster.
-A cluster extends into the directions with $\Upsilon_{k}>0$ while it shrinks in the others for half period of oscillation, with particles reversing their direction of motion during the remaining half period.
-In the following, we assume $\Upsilon_{1}<0$ and order the eigenvalues so that $|\Upsilon_{i}|>|\Upsilon_{i+1}|$.
-In two dimensions, the traceless condition implies $\Upsilon_{1}=-\Upsilon_{2}$.
-Figure 4(b) illustrates the quadrupoles associated with each cluster on a heat map of the magnitude of the coarse-grained cage-relative displacement, finding expected correlations in light of panel (a).
-The figure illustrates a small fraction of the system and represents each quadrupole with four arrows with a length proportional to ±$\pm\Upsilon_{i}{\bm{U}}_{i}$.
-In three dimensions, we examine the relevance of various deformation modes that satisfy the traceless condition $\sum\Upsilon_{i}=0$ by analyzing the probability distribution $P(\Upsilon_{1},\Upsilon_{2})$.
-This distribution is supported within a range defined by the line $\Upsilon_{1}+\Upsilon_{2}=0$, which represents pure shear deformation, and the line $\Upsilon_{1}+2\Upsilon_{2}=0$, which corresponds to volume-preserving isotropic ($\Upsilon_{3}=\Upsilon_{2}=-\Upsilon_{1}/2$) uniaxial compressions.
-Figure 4(c) shows that quadrupolar deformations are predominantly pure shear, as $P(\Upsilon_{1},\Upsilon_{2})$ peaks near the $\Upsilon_{1}+\Upsilon_{2}=0$ line for different values of $x_{c}$.
-This result is further supported by direct visualization of the quadrupoles (Figure 4(d)), depicted as collections of 6 arrows with lengths proportional to ±$\pm\Upsilon_{i}{\bm{U}}_{i}$.
+2(d) presents density maps of the $P(\sigma_{1},\sigma_{2})$ probability distribution at $x_{c}=1.5$ and $x_{c}=1.7$ (shifted vertically) and suggests that the localised defects are three-dimensional compact (panel b) clusters with a high degree of anisotropy.
+Due to this anisotropy, the radius of gyration scales as $R_{g}\propto n^{q}$, with $q\simeq 0.55$ in 2D and $q\simeq 0.5$ in 3D, if not for large clusters, as shown in Fig.
+2(e).
+The distribution of the clusters’ radius of gyration is exponentially distributed (see Appendix D), allowing us to associate a typical length scale $\lambda$ to each $x_{c}$.
+Fig.
+2(f) shows that $\omega_{\rm bp}\propto c_{s}/\lambda$, in both two and three dimensions, proving these defects’ relevance to the boson peak phenomenology.
+Figure 3: 1 Heat map of the magnitude of the coarse-grained cage-relative displacement, normalised by its average value, with superimposed quadrupolar displacements of the defects.
+The figure refers to $x_{c}=1.7$ and depicts 1/12th of the system’s area.
+b Density map of the two largest (in modulus) eigenvalues of the deviatoric strain associated with the defects, normalised to its maximum value.
+The density is higher close to the $\Upsilon_{2}=-\Upsilon_{1}$ line, implying that the quadrupoles correspond to pure shear deformations.
+c We represent the quadrupolar deformations by attaching to each defect’s centre of mass the displacements along the principal axis $\pm\Upsilon_{1}{\bm{u}}_{1}$ (blue), $\pm\Upsilon_{2}{\bm{u}}_{2}$ (green) and $\pm\Upsilon_{3}{\bm{u}}_{3}$ (red).
+Since quadrupoles mostly correspond to pure-shear deformations, ${\bf u}_{3}$ is barely visible.
+Vibrational properties – To investigate the vibrational properties of these clusters, we first evaluate the strain tensors coarse-grained over a small length scale.
+Since the strain is insensitive to local uniform translations, as the cage-relative measure, the cage-relative clusters occur in regions of high local strain.
+This allows us to evaluate a defect strain tensor $\mathcal{E}=\mathcal{E}_{\rm vol}+\mathcal{E}_{\rm dev}$, as detailed in Appendix E.
+We find that these tensors are dominated by their deviatoric component, indicating that defects perform volume-preserving deformations, as $R=\frac{\sqrt{Q}}{|\rm{Tr}(\mathcal{E})|}\gg 1$.
+Here, $Q=1/2\sum\Upsilon_{k}^{2}$ is the quadrupolar charge and $\Upsilon_{k}$ are the eigenvalues of the traceless $\mathcal{E}_{\rm dev}$, and $|\rm{Tr}(\mathcal{E})|$ quantifies the degree of volumetric deformations.
+In the following, we assume $\Upsilon_{1}<0$, order the eigenvalues so that $|\Upsilon_{i}|>|\Upsilon_{i+1}|$, and indicate their associated eigenvectors with ${\bf U}_{k}$.
+A cluster extends into the directions with $\Upsilon_{k}>0$ while it shrinks in the others for half a period of oscillation, with particles reversing their direction of motion during the remaining half period.
+The traceless condition implies $\Upsilon_{1}=-\Upsilon_{2}$ in two dimensions.
+Figure 3(a) illustrates the quadrupoles associated with each cluster on a heat map of the magnitude of the coarse-grained cage-relative displacement, finding expected correlations.
+The figure illustrates a small fraction of the system and represents each quadrupole with four arrows with a length proportional to $\pm\Upsilon_{i}{\bm{U}}_{i}$.
+See Fig.
+5 in Appendix A for a snapshot of the whole system and illustrations of the underlying displacement field.
+In three dimensions, different deformation modes satisfy the traceless condition $\sum\Upsilon_{i}=0$.
+We expose the dominant deformation mode by analyzing the probability distribution $P(\Upsilon_{1},\Upsilon_{2})$.
+In the $\Upsilon_{1}$-$\Upsilon_{2}$ plane, this distribution is supported within a region bounded by the line $\Upsilon_{1}+\Upsilon_{2}=0$, which represents pure shear deformation, and the line $\Upsilon_{1}+2\Upsilon_{2}=0$, which corresponds to volume-preserving isotropic ($\Upsilon_{3}=\Upsilon_{2}=-\Upsilon_{1}/2$) uniaxial compressions.
+Figure 3(b) shows that quadrupolar deformations are predominantly pure shear, as $P(\Upsilon_{1},\Upsilon_{2})$ peaks near the $\Upsilon_{1}+\Upsilon_{2}=0$ line for different values of $x_{c}$.
+This result is further supported by direct visualization of the quadrupoles (Figure 3(c)), depicted as collections of 6 arrows with lengths proportional to $\pm\Upsilon_{i}{\bm{U}}_{i}$.
 Since $|\Upsilon_{3}|\ll|\Upsilon_{2}|$, the blue arrows corresponding to displacements in the $\bm{U}_{3}$} direction are not visible, making the quadrupoles appear as a two-dimensional cross.
-Figure 4: Quadrupolar nature of defects.
-a Pearson correlation coefficient between coarse-grained cage-relative displacement and quadrupolar charge.
-b Quadrupoles associated with the cluster superimposed on a heat map of the magnitude of the coarse-grained cage-relative displacement, normalised by its average value.
-The figure refers to $x_{c}=1.7$ and depicts 1/12th of the system’s area.
-c Density map of the two largest (in modulus) eigenvalues of the deviatoric strain associated with the defects, normalised to its maximum value.
-The density is higher close to the $\Upsilon_{2}=-\Upsilon_{1}$ line, implying that the quadrupoles correspond to pure shear deformations.
-d We represent the quadrupolar deformations by attaching to each defect’s centre of mass the displacements along the principal axis ±$\pm\Upsilon_{1}{\bm{u}}_{1}$ (blue), ±$\pm\Upsilon_{2}{\bm{u}}_{2}$ (green) and ±$\pm\Upsilon_{3}{\bm{u}}_{3}$ (red).
-Since quadrupoles mostly correspond to pure-shear deformations, ${\bf u}_{3}$ is barely visible.
-The localised defects regulate the boson peak
-We test more rigorously whether quadrupolar defects regulate the boson peak by evaluating if their density $\rho_{d}$ (number per unit volume) accounts for the excess density of vibrational modes defined as $\Delta D(\omega_{\rm bp})=D(\omega_{\rm bp})-D_{\rm Debye}(\omega_{\rm bp})$, where $D_{\rm Debye}(\omega_{\rm bp})=A_{D}\omega^{D-1}$ represents Debye’s prediction, as expected if the phonon-defect interaction is weak.
-In making this test, it is important to note that $\rho_{d}$ is not uniquely determined, as it depends on the definition of defects used.
-We define defects as clusters of particles with cage-relative displacements in the top 5%.
-This threshold is not critical, as the SI shows that varying it does not significantly affect the number of defects or $\rho_{d}$, as it primarily impacts the defect size.
-Instead, the density of defects is more influenced by the minimum number of particles, $n_{\rm min}$, required for a cluster to be considered a defect.
-This threshold is introduced to ensure that volume-preserving deformations involve more than just a few particles.
-Fig.
-5(a) explores the dependence of the density of defects on $x_{c}$ in two-dimensional systems.
-An $n_{\rm min}$ increases, $\rho_{d}$ quickly looses any $x_{c}$ dependence.
-Similarly, we show in panel (b) that the excess of modes $\Delta D(\omega_{\rm bp})$ is also constant in two dimensions, showing therefore a clear correlation with the also constant $\rho_{d}$.
-In three dimensions, Fig.
-5(c) demonstrates that the density of defects increases as $x_{c}$ decreases, consistent with the loss of stability.
-The excess of modes follows the same trend with $x_{c}$, with Fig.
-5(d) demonstrating their proportionality.
-Importantly, in 2D, $\Delta D(\omega_{\rm bp})$ and $\rho_{d}$ have comparable values, besides being both $x_{c}$ independent.
-In 3D, we find that $\Delta D(\omega_{\rm bp})=c\rho_{d}$.
-The constant of proportionality $c$ depends on the specific definition of defects, but it is consistently of order 1.
-These findings strongly suggests that the excess of vibrational modes at the boson peak frequency originates from the quadrupolar defects we have identified.
-Figure 5: Density of vibrational defects and excess density of modes.
-We evaluated the excess density of vibrational modes at the boson peak frequency, $\Delta D(\omega_{\rm bp})=D(\omega_{\rm bp})-D_{\rm Debye}(\omega_{\rm bp})$, $D_{\rm Debye}(\omega)$ being Debye’s density of state, and the density $\rho_{d}$ of vibrational defects containing more than $n$ particles.
-In two dimensions, a $\rho_{d}$ for clusters containing a minimum number of particle $n_{\rm min}$ is $x_{c}$ independent.
-Similarly, b, $\Delta D(\omega_{\rm bp})$ is $x_{c}$ independent.
-In three dimensions, c $\Delta D(\omega_{\rm bp})$ grows with $x_{c}$, and d $\rho_{d}\propto\Delta D(\omega_{\rm bp})$.
-Discussion and relation to stringlets
-We have identified localised defects hybridizing with plane waves at the boson peak frequency through a filtering technique based on the physical insight that localized vibrations are generally less affine than extended ones.
-This technique allowed us to isolate the vibrational defects and study their geometrical and vibrational properties.
-The vibrational defects have some degree of spatial anisotropy but are compact rather than planar or string-like.
-Their oscillatory motion is quadrupolar.
-Specifically, the defects undergo pure shear oscillatory deformations, extending along one axis while contracting along a perpendicular one in both two and three spatial dimensions.
-Analysis of various two- and three-dimensional systems shows that the number density of these defects quantitatively accounts for the excess of modes at the boson peak, confirming their physical relevance.
-At this point, one might then wonder how our findings are compatible with the recent claims by Hu and Tanaka [25], that on the contrary identify the microscopic origin of the BP in one-dimensional string-like dynamical objects.
-These ‘stringlets‘ are defined as clusters of particles with a large vibrational amplitude, or equivalently, a high ‘vibrability’ $\Psi_{i}=\sum_{k}\frac{1}{\omega_{k}^{2}}\left|{\bf e}_{k,i}\right|^{2}$, where the sum extends over all modes with frequencies $\omega_{k}$ close to the boson peak frequency $|\omega_{\text{bp}}-\omega_{k}|<\Delta\omega$.
-We have performed the same stringlet analysis in our simulation model.
-As shown in Figure 6(a) by the average coordination number of the particles participating in the stringlets, we find that the clusters defined as in [25] are compact and extended objects, rather than one-dimensional, very similar to the localised defects found in our work (Figure 3(b)).
-In Figure 6(b), we observe a high spatial correlation between particles with vibrability in the top 5% (red) alongside those with cage-relative displacements in the top 5% (blue), indicating a high correlation between vibrational motion and cage-relative displacement at the BP frequency.
-These correlations are further confirmed by measuring the fraction $f$ of particles with high vibrability and cage-relative displacement, which would be $f_{\text{rand}}=0.05$ in the absence of correlations.
-Figure 6(c) shows that $f$ is independent of $x_{c}$ and significantly exceeds the random expectation.
-Our analysis provides strong evidence that quadrupolar localised defects and high-vibrability clusters (previously identified as stringlets) are two sides of the same coin.
-This reconciles two of the most successful theories for the origin of the BP in glasses and reveal the ultimate nature of the defects responsible for it.
-Along the lines of [15], it would be interesting to understand whether the same defective structures could be also rationalized by means of shear modulus fluctuations and whether our findings are compatible with another successful theory of the BP – heterogeneous elasticity theory [45].
-Moreover, it would be fruitful to investigate the relation between the defects discussed in our manuscript and (I) the geometric charges responsible for plastic screening and anomalous elastic response [46], (II) the topological vortex-like defects recently connected to plastic soft spots [47, 48].
-Figure 6: Reconciling stringlets and quadrupolar defects.
-a The coordination number of vibrability clusters increases with $n$, following a trend similar to that of clusters formed based on cage-relative displacement (Fig.
-3c).
+Overall, these investigations show that modes at the boson peak comprise plane waves hybridized with two-dimensional defects that perform pure shear deformations, in both two- and three-spatial dimensions.
+Figure 4: a The coordination number of vibrability clusters increases with $n$, following a trend similar to that of clusters formed based on cage-relative displacement (see Fig.3b).
 b Particles exhibiting high cage-relative displacement at the boson peak (red) and high vibrability (blue), both in the top 5%, for $x_{c}=1.7$.
@@
 c The fraction of particles displaying both large cage-relative displacement and high vibrability significantly exceeds the expectation under random conditions.
-Methods
-Numerical model
+The connection with ‘stringlets’ – At this point, one might then wonder how our findings are compatible with the recent claims by Hu and Tanaka [28], who suggested that the BP is induced by one-dimensional clusters, or ‘stringlets’.
+These are clusters of particles with a large vibrational amplitude, or equivalently, a high ‘vibrability’ $\Psi_{i}=\sum_{k}\frac{1}{\omega_{k}^{2}}\left|{\bf e}_{k,i}\right|^{2}$, where the sum extends over all modes with frequencies $\omega_{k}$ close to the boson peak frequency.
+We have repeated their analysis in our simulation model.
+Figure 4(a) shows that the average coordination number of the particles participating in the resulting cluster is large, indicating that these are compact and extended objects, rather than one-dimensional, very similar to the localised defects investigated in Fig.
+2(b).
+In Figure 4(b), we observe a high spatial correlation between particles with vibrability in the top 5% (blue) alongside those with cage-relative displacements in the top 5% (red), indicating a high correlation between vibrational motion and cage-relative displacement at the BP frequency.
+These correlations are further confirmed by measuring the fraction $f$ of particles with high vibrability and cage-relative displacement, which is much larger than expected in the absence of correlations, as we illustrate in Figure 4(c).
+Overall, our analysis provides strong evidence that quadrupolar localised defects and high-vibrability clusters are two sides of the same coin.
+In our systems, however, these clusters are not string-like, and their motion is quadrupolar rather than one-dimensional.
+Conclusions – These findings support earlier results [18, 19] suggesting a relation between the boson peak and quasi-localised vibrational modes in a soft-potential model perspective [6, 7, 8, 9].
+Our approach has the merit of allowing for a straightforward identification of the hybridized defects in 2D, when these are not visible in the displacement field, and in 3D.
+The connection between our defects and QLMs deserves further investigation.
+In particular, while our defects exclusively undergo pure shear deformations, Ref.
+[49] reported a low-frequency QLM undergoing simple shear deformations where extension occurs in one direction and contraction in the orthogonal ones.
+In addition, we find that the displacement is predominantly two-dimensional, even in three dimensions, and hence, it does not align with that of the low-frequency localized modes [50].
+Along the lines of [18], it would be interesting to understand whether the same defective structures could also be rationalized using shear modulus fluctuations and whether our findings are compatible with another successful theory of the BP – heterogeneous elasticity theory [51].
+Moreover, it would be fruitful to investigate the relation between the defects discussed in our manuscript and (i) the geometric charges responsible for plastic screening and anomalous elastic response [52], (ii) the topological vortex-like defects recently connected to plastic soft spots [53, 54].
+Acknowledgements- MPC acknowledges support by the Singapore Ministry of Education through grants MOE-T2EP50221-0016 and RG152/23.
+CJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01).
+MB acknowledges the support of the sponsorship from the Yangyang Development Fund.
+Appendix A Extracting defects from a disordered vibrational mode
+Fig.
+5a presents an eigenmode at the boson peak frequency, characterized by swirls that indicate hybridization with plane waves.
+This hybridization prevents visually isolating quadrupolar-like defects as recently proposed [19, 20].
+Fig.
+5b displays the defects identified by our proposed filtering technique, while panel c provides a magnified view of a region covering 1/12th of the system’s area, offering a clearer visualization of the quadrupolar-like localized defects.
+Figure 5: a Eigenvector at the boson peak frequency for a two-dimensional system with $x_{c}=1.7$.
+It is arduous to spot in similar modes isolated quadrupolar-looking defects [19, 20].
+b Defects identified by our approach.
+c A zoom on a region corresponding to 1/12th of the system’s area.
+Appendix B Numerical Model
 We consider two- and three-dimensional systems of particles of diameter $\sigma_{i}$, drawn from a uniform random distribution in the range [0.8:1.2], that interact via family of LJ-like potentials consisting of a repulsive and an attractive part, $U(r_{ij},x_{c})=U_{r}(r_{ij})+U_{a}(r_{ij},x_{c})$ [34, 35].
@@
 The attractive part, which only acts for distances in the range $2^{1/6}\sigma_{ij}\leq r_{ij}\leq x_{c}\sigma_{ij}$, is given by
-$U_{a}(r_{ij})=\epsilon\left[a_{0}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{12}-a_{1}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}+\sum_{l=0}^{3}{c_{2l}\left(\frac{r_{ij}}{\sigma_{ij}}\right)^{2l}}\right].$ (2)
+$U_{a}(r_{ij})=\epsilon\left[a_{0}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{12}\hskip-11.38092pt-a_{1}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}\hskip-5.69046pt+\sum_{l=0}^{3}{c_{2l}\left(\frac{r_{ij}}{\sigma_{ij}}\right)^{2l}}\right].$ (2)
 The parameters $a_{0},a_{1}$ and $c_{2l}$ are chosen such that the potential $U(r_{ij})$ and its first two derivatives are continuous at the minimum $r_{ij}^{\rm min}$ and at the cutoff $r_{ij}^{(c)}=x_{c}\sigma_{ij}$, where the potential also vanishes.
@@
 This parameter influences the relaxation dynamics [35] and the mechanical response [34, 36, 37, 38].
-Data availability
-The datasets generated and analysed during the current study are available upon reasonable request by contacting the corresponding authors.
-Funding
-MPC acknowledges support by the Singapore Ministry of Education through grants MOE-T2EP50221-0016 and RG152/23.
-CJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01).
-MB acknowledges the support of the sponsorship from the Yangyang Development Fund.
-Author contributions statement
-M.P.C.
-and M.B.
-devised the project.
-M.P.C., S.M.
-and M.B.
-wrote the paper.
-S.M., D.S.Y.H.
-performed the numerical simulations and prepared the figures.
-All authors contributed to the scientific discussion and refinement of the manuscript.
-Supplementary information
-Density of vibrational modes
-We compare the density of vibrational modes obtained from Fourier transforming the velocity autocorrelation function (points) and via the direct diagonalisation of the Hessian matrix in Fig.
-7 for systems with 256k particles.
-In the figure, we also illustrate the Debye’s prediction, $D_{\rm Debye}=A_{D}\omega^{D-1}$.
-In 2D, $A_{D}=\frac{2}{8\pi\rho/(c_{l}^{-2}+c_{s}^{-2})}$, with $\rho$ the number density and $c_{l}=\sqrt{\frac{K+\mu}{\rho}}$ and $c_{s}=\sqrt{\frac{\mu}{\rho}}$ the longitudinal and the transverse shear velocities.
-In 3D, $A_{D}=\frac{3}{18\pi^{2}\rho/(c_{l}^{-3}+2c_{s}^{-3})}$, $c_{l}=\sqrt{\frac{K+\frac{4\mu}{3}}{\rho}}$ and $c_{s}=\sqrt{\frac{\mu}{\rho}}$.
-We determine the bulk $K$ and shear $\mu$ moduli by investigating the stress-strain relationships in the linear response regime.
-Figure 7: Different approaches to evaluate the vDOS.
-We compare diverse approaches to evaluating the vDOS, in 2D and 3D.
-Evaluating the vDOS by Fourier transforming the velocity autocorrelation function and via the diagonalization of the Hessian gives consistent results.
-Due to our system’s large size, 256k particles, we compute only the lower end of the spectrum via the direct diagonalization of the Hessian.
-Cage-relative measures filter plane waves at the boson peak
+We evaluate the vDOS over a broad frequency range by computing the Fourier transform of the velocity autocorrelation function in the linear response regime by linearizing the equation of motion, implying the absence of anharmonic effects.
+Additionally, we obtain the vDOS by diagonalizing the Hessian matrix, which provides access to eigenvectors, ${\bf e}_{k}$, and eigenfrequencies, $\omega_{k}$, in the low-frequency regime, including the boson peak.
+These methods yield consistent results.
+Fig.6(a) shows the reduced vDOS, $\color[rgb]{0,0,0}D(\omega)/A_{D}\omega^{d-1}$}, with the Debye prediction, $\color[rgb]{0,0,0}D_{\rm Debye}=A_{D}\omega^{d-1}$}, highlighting the BP in two and three spatial dimensions.
+Here, $A_{D}=d/\omega_{D}^{d}$ where $d$ is dimension, $\omega_{D}^{d}=\frac{18\pi^{2}\rho}{c_{l}^{-3}+2c_{s}^{-3}}$ in 3d and $\omega_{D}^{d}=\frac{8\pi\rho}{c_{l}^{-2}+c_{s}^{-2}}$ in 2d, with the longitudinal and shear velocities given by $c_{l}=\sqrt{\frac{K+\frac{4}{3}\mu}{\rho}}$ in 3d, $c_{l}=\sqrt{\frac{K+\mu}{\rho}}$ in 2d and $c_{s}=\sqrt{\frac{\mu}{\rho}}$ in both 2d and 3d.
+Here, $K$ is the bulk modulus and $\mu$ is the shear modulus, which we measure by investigating the stress-strain relationships in the linear response regime.
+In 3D, increasing $x_{c}$ raises the peak amplitude (see Fig.
+6b) while lowering its frequency, $\omega_{\rm bp}$ (panel c), suggesting that shorter attractive wells stabilize glasses, as previously proposed [18, 36, 38, 34].
+In 2D, the effect of $x_{c}$ is weaker, indicating that the material stability, which depends on $x_{c}$ in our model, has a dimensionality-dependent influence on the BP.
+It remains to be ascertained whether these results are universal across different systems.
+For 3D, we analyze $x_{c}\geq 1.5\sigma$, as smaller values maintain a phonon band structure at high frequencies, complicating $\omega_{\rm bp}$ identification, as illustrated in Fig.
+7.
+In 2D, we consider $x_{c}\geq 1.4\sigma$, since for lower values the BP broadens, making $\omega_{\rm bp}$ difficult to estimate.
+Figure 6: a The reduced vDOS in three and (inset) two dimensions for systems differing in the extension $x_{c}$ of the intermolecular attractive potential well.
+In 3D, the materials lose stability as $x_{c}$ increases, as the boson peak intensity increases (b) while its frequency decreases (c).
+In 2D, attraction does not sensibly influence the excess vibrational modes.
+Frequencies are expressed in units of the natural frequency $\omega_{0}=c_{s}/a_{0}$, with $a_{0}$ the interparticle spacing and $c_{s}$ the shear wave speed.
+The inset of (c) illustrates the localised defects that, according to our investigation, induce the boson peak in a three-dimensional system with $x_{c}=1.7$.
+Figure 7: Density of states normalized by Debye’s prediction for $x_{c}=1.4$ in 3d.
+In finite systems, at low frequencies, the discrete nature of the spectrum emerges, and hence, $D(\omega)/A_{D}\omega^{2}$ approaches zero rather than a constant.
+At small $x_{c}$, the discrete nature of the spectrum extends to higher frequencies, making it difficult to identify the boson peak frequency.
+Appendix C Cage-relative filtering approach
+To prove that our filtering approach does not filter out localized vibrations, we illustrate in Fig.
+8(a) a snapshot of a two-dimensional system where each particle is coloured based on its $|{\bm{e}}_{k,i}|^{2}$, for a low-frequency localised mode $k$.
+The localised excitation remains visible in panel (b), where particles are coloured according to the magnitude of their cage-relative displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$, proving that our approach preserves localized defects.
+Panels (c) and (d) illustrate the same results for a hybridized plane wave.
+Figure 8: a Colour map of the magnitude $|{\bm{e}}_{k,i}|^{2}$ of the particle displacement associated with a low-frequency mode with $\omega_{k}=0.23\omega_{\rm bp}$.
+A localized defect is visible.
+b The defect remains visible in the colour map of the magnitude of the cage-relative displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$.
+In all plots, the magnitudes are scaled so that their average value is 11.
+c,d report analogous results at a similar frequency for a hybridized rather than a quasi-localized mode.
+Figure 9: Squared norm of the cage-relative displacement field $|{\bm{p}}^{\rm cr}|^{2}$ associated with phonons ($|{\bm{p}}|^{2}=1$) of wavelength $\lambda$, in two dimensions.
+The vertical lines mark the wavelength $\lambda_{\rm bp}$ of phonons at the boson peak frequency for the extremal investigated $x_{c}$ values.
+Analogous results hold in 3D.
 To evaluate the cage-relative ability to filter extended modes of norm $|{\bm{u}}|^{2}=1$, we consider $|{\bm{u}}^{\rm cr}|^{2}$.
@@
 There is no need to consider the superposition of phonons as the cage-relative ‘operator’ is linear, $(a{\bm{u}}+b{\bm{v}})^{\rm cr}=a{\bm{u}}^{\rm cr}+b{\bm{v}}^{\rm cr}$.
-Any linear combination of ‘filtered’ displacement fields is also filtered.
 Fig.
 9 illustrates $|{\bm{p}}^{\rm cr}|^{2}$ for phonon $\bf p$} of wavelength $\lambda$, in two dimensions.
-The larger the phonon wavelength, the more efficient the filtering.
+Filtering is more efficient at larger wavelengths, where the induced deformation is more affine.
 Importantly, filtering is highly efficient at the wavelength $\lambda_{\rm bp}=\frac{2\pi c_{s}}{\omega_{\rm bp}}$ of the phonons expected to hybridize at the boson peak, as $|{\bm{p}}^{\rm cr}|^{2}$ is in the range $10^{-4}$ to $10^{-3}$, its value depending on the cutoff $x_{c}$ of the interaction potential.
-Figure 8: Filtering phonons via the cage-relative measured.
-Squared norm of the cage-relative displacement field $|{\bm{p}}^{\rm cr}|^{2}$ associated with phonons ($|{\bm{p}}|^{2}=1$) of wavelength $\lambda$, in two dimensions.
-The vertical lines mark the wavelength $\lambda_{\rm bp}$ of phonons at the boson peak frequency for the extremal investigated $x_{c}$ values.
-Analogous results hold in 3D.
-Density of vibrational defects
-We define defects as clusters of particles whose cage-relative displacement is within the top $c=5\%$.
-We explore how the density of defects, $\rho_{d}$, depends on the percentage $c$ of particles with the highest cage-relative displacement.
-As $c\to 0$, $\rho_{d}$ vanishes since there are no clusters, meaning no particles meet the clustering criteria.
-As $c$ increases, $\rho_{d}$ grows, reflecting both the increasing size and number of clusters.
-However, beyond a certain point, as clusters begin to merge, $\rho_{d}$ decreases, leading to a maximum in defect density at an intermediate $c$ value.
+Appendix D Cluster linear size
 Fig.
-6a shows that, in 2D, the density of defects remains $x_{c}$ independent for $c<10\%$.
-Unless this threshold is considered unrealistically high, the exact value of $c$ does not significantly affect our conclusions.
-In 3D, the behavior of $\rho_{d}$ for small $x_{c}$ values is illustrated in Fig.
-6c: $\rho_{d}$ remains unchanged between $x_{c}=1.8$ and $x_{c}=1.7$ and increases as $x_{c}$ decreases further.
-Thus, our results are not highly sensitive to the exact choice of $x_{c}$.
-Figure 9: Cutoff dependence of the density of defects.
-We perform a cluster analysis of the cof the particles with the largest magnitude of the cage relative displacement (in the mode at boson peak frequency) to identify the defects.
-a the density of defects $\rho_{d}$ as a maximum as a function of $c$, and is roughly constant as $c$ varies from $5\%$ to $7\%$.
-b $\rho_{d}$ remains constant in this $c$ range for all $x_{c}$, also when we consider as defects only cluster with at least $n_{\rm min}$ particles.
-We illustrate the considered values of $n_{\rm min}$ for $x_{c}=1.5$.
+10 shows that the radius of gyration of the clusters is exponentially distributed, at large $R_{g}$, $P(R_{g})=ae^{-R_{g}/\lambda}$, which allow us to associate a typical length scale $\lambda$ to each system.
+Figure 10: Distribution of the defects’ radius of gyration, in 2D and 3D, for diverse values of $x_{c}$.
+Appendix E Cluster deviatoric strain
+We evaluate the local strain tensor ${\bm{\epsilon}}=\frac{1}{2}\left({\bm{\nabla}}{\bm{e}}+({\bm{\nabla}}\bm{e})^{T}\right),$ associated with the displacement field (eigenvector, $\bm{e}$}) of the mode at the boson peak frequency, coarse-grained on grids of side-length $w\simeq 1.1$.
+We decompose this tensor into its volumetric and deviatoric components: $\bm{\varepsilon}=\bm{\varepsilon}_{\rm vol}+\bm{\varepsilon}_{\rm dev}$.
+Here, $\bm{\varepsilon}_{\rm vol}$ represents local volumetric deformations.
+We define a cluster deviatoric strain as $\mathcal{E}_{\rm dev}=\sum\bm{\varepsilon}_{\rm dev}$, where the sum extends over all grid cells containing particles of the considered cluster.